Classify each of the species given below as parama
Answer #1
Concepts and reason

The given problem is based on the concept of Molecular Orbital Theory (MOT). The MOT can be used to predict whether the molecule is paramagnetic or diamagnetic.

If all the electrons are paired, the molecule is diamagnetic and if the molecule has one or more unpaired electrons, it is paramagnetic.

Fundamentals

To classify the molecules as paramagnetic or diamagnetic, write the molecular orbital configurations of the molecules based on MOT. Then classify on the basis of paired and unpaired electrons.

The molecular orbital configuration of O2{{\rm{O}}_{\rm{2}}} is as follows:

σ1s2σ1s2σ2s2σ2s2σ2px2{π2py2,π2pz2}{π2py1,π2pz1}\sigma 1{s^2}{\sigma ^*}1{s^2}\sigma 2{s^2}\sigma {}^*2{s^2}\sigma 2p_x^2\left\{ {\pi 2p_y^2,\pi 2p_z^2} \right\}\left\{ {{\pi ^*}2p_y^1,{\pi ^*}2p_z^1} \right\}

The molecular orbital configuration of O22{\rm{O}}_2^{2 - } is as follows:

σ1s2σ1s2σ2s2σ2s2σ2px2{π2py2,π2pz2}{π2py2,π2pz2}\sigma 1{s^2}{\sigma ^*}1{s^2}\sigma 2{s^2}\sigma {}^*2{s^2}\sigma 2p_x^2\left\{ {\pi 2p_y^2,\pi 2p_z^2} \right\}\left\{ {{\pi ^*}2p_y^2,{\pi ^*}2p_z^2} \right\}

The molecular orbital configuration of O22+{\rm{O}}_2^{2 + } is as follows:

σ1s2σ1s2σ2s2σ2s2σ2px2{π2py2,π2pz2}\sigma 1{s^2}{\sigma ^*}1{s^2}\sigma 2{s^2}\sigma {}^*2{s^2}\sigma 2p_x^2\left\{ {\pi 2p_y^2,\pi 2p_z^2} \right\}

Hence, O2{{\rm{O}}_{\rm{2}}} is paramagnetic as it has two unpaired electrons. O22{\rm{O}}_2^{2 - } and O22+{\rm{O}}_2^{2 + } are diamagnetic as all the electrons are paired.

The molecular orbital configuration of F2 is:

σ1s2σ1s2σ2s2σ2s2σ2px2{π2py2,π2pz2}{π2py2,π2pz2}\sigma 1{s^2}{\sigma ^*}1{s^2}\sigma 2{s^2}\sigma {}^*2{s^2}\sigma 2p_x^2\left\{ {\pi 2p_y^2,\pi 2p_z^2} \right\}\left\{ {{\pi ^*}2p_y^2,{\pi ^*}2p_z^2} \right\}

The molecular orbital configuration of F22+{\rm{F}}_2^{2 + } is:

σ1s2σ1s2σ2s2σ2s2σ2px2{π2py2,π2pz2}{π2py1,π2pz1}\sigma 1{s^2}{\sigma ^*}1{s^2}\sigma 2{s^2}\sigma {}^*2{s^2}\sigma 2p_x^2\left\{ {\pi 2p_y^2,\pi 2p_z^2} \right\}\left\{ {{\pi ^*}2p_y^1,{\pi ^*}2p_z^1} \right\}

The molecular orbital configuration of F22{\rm{F}}_2^{2 - } is:

σ1s2σ1s2σ2s2σ2s2σ2px2{π2py2,π2pz2}{π2py2,π2pz2}σ2px2\sigma 1{s^2}{\sigma ^*}1{s^2}\sigma 2{s^2}\sigma {}^*2{s^2}\sigma 2p_x^2\left\{ {\pi 2p_y^2,\pi 2p_z^2} \right\}\left\{ {{\pi ^*}2p_y^2,{\pi ^*}2p_z^2} \right\}{\sigma ^*}2p_x^2

Hence, F2 and F22{\rm{F}}_2^{2 - } are diamagnetic and F22+{\rm{F}}_2^{2 + } is paramagnetic.

The molecular orbital configuration of N2{{\rm{N}}_2} is as follows:

σ1s2σ1s2σ2s2σ2s2{π2py2,π2pz2}σ2px2\sigma 1{s^2}{\sigma ^*}1{s^2}\sigma 2{s^2}\sigma {}^*2{s^2}\left\{ {\pi 2p_y^2,\pi 2p_z^2} \right\}\sigma 2p_x^2

The molecular orbital configuration of N2+{\rm{N}}_{\rm{2}}^{\rm{ + }} is as follows:

σ1s2σ1s2σ2s2σ2s2{π2py2,π2pz2}σ2px1\sigma 1{s^2}{\sigma ^*}1{s^2}\sigma 2{s^2}\sigma {}^*2{s^2}\left\{ {\pi 2p_y^2,\pi 2p_z^2} \right\}\sigma 2p_x^1

The molecular orbital configuration of N2{\rm{N}}_{\rm{2}}^ - is as follows:

σ1s2σ1s2σ2s2σ2s2{π2py2,π2pz2}σ2px2{π2py1,π2pz0}\sigma 1{s^2}{\sigma ^*}1{s^2}\sigma 2{s^2}\sigma {}^*2{s^2}\left\{ {\pi 2p_y^2,\pi 2p_z^2} \right\}\sigma 2p_x^2\left\{ {{\pi ^*}2p_y^1,{\pi ^*}2p_z^0} \right\}

The molecular orbital configuration of N22+{\rm{N}}_{\rm{2}}^{{\rm{2 + }}} is as follows:

σ1s2σ1s2σ2s2σ2s2{π2py2,π2pz2}\sigma 1{s^2}{\sigma ^*}1{s^2}\sigma 2{s^2}\sigma {}^*2{s^2}\left\{ {\pi 2p_y^2,\pi 2p_z^2} \right\}

Hence, N2{{\rm{N}}_2} and N22+{\rm{N}}_2^{2 + } are diamagnetic while N2+{\rm{N}}_{\rm{2}}^{\rm{ + }} and N2{\rm{N}}_{\rm{2}}^ - are paramagnetic.

The molecular orbital configuration of C2{{\rm{C}}_2} is as follows:

σ1s2σ1s2σ2s2σ2s2{π2py2,π2pz2}\sigma 1{s^2}{\sigma ^*}1{s^2}\sigma 2{s^2}\sigma {}^*2{s^2}\left\{ {\pi 2p_y^2,\pi 2p_z^2} \right\}

The molecular orbital configuration of C22+{\rm{C}}_{\rm{2}}^{{\rm{2 + }}} is as follows:

σ1s2σ1s2σ2s2σ2s2{π2py1,π2pz1}\sigma 1{s^2}{\sigma ^*}1{s^2}\sigma 2{s^2}\sigma {}^*2{s^2}\left\{ {\pi 2p_y^1,\pi 2p_z^1} \right\}

The molecular orbital configuration of C22{\rm{C}}_{\rm{2}}^{2 - } is as follows:

σ1s2σ1s2σ2s2σ2s2{π2py2,π2pz2}σ2px2\sigma 1{s^2}{\sigma ^*}1{s^2}\sigma 2{s^2}\sigma {}^*2{s^2}\left\{ {\pi 2p_y^2,\pi 2p_z^2} \right\}\sigma 2p_x^2

Hence, C2{{\rm{C}}_2} and C22{\rm{C}}_{\rm{2}}^{{\rm{2 - }}} are diamagnetic while C22+{\rm{C}}_{\rm{2}}^{{\rm{2 + }}} is paramagnetic.

The molecular orbital configuration of B2{{\rm{B}}_2} is as follows:

σ1s2σ1s2σ2s2σ2s2{π2py1,π2pz1}\sigma 1{s^2}{\sigma ^*}1{s^2}\sigma 2{s^2}\sigma {}^*2{s^2}\left\{ {\pi 2p_y^1,\pi 2p_z^1} \right\}

The molecular orbital configuration of B22+{\rm{B}}_{\rm{2}}^{{\rm{2 + }}} is as follows:

σ1s2σ1s2σ2s2σ2s2\sigma 1{s^2}{\sigma ^*}1{s^2}\sigma 2{s^2}\sigma {}^*2{s^2}

The molecular orbital configuration of B22{\rm{B}}_{\rm{2}}^{2 - } is as follows:

σ1s2σ1s2σ2s2σ2s2{π2py2,π2pz2}\sigma 1{s^2}{\sigma ^*}1{s^2}\sigma 2{s^2}\sigma {}^*2{s^2}\left\{ {\pi 2p_y^2,\pi 2p_z^2} \right\}

Hence, B2{{\rm{B}}_2} is paramagnetic while B22+{\rm{B}}_{\rm{2}}^{{\rm{2 + }}} and B22{\rm{B}}_{\rm{2}}^{2 - } are diamagnetic.

The molecular orbital configuration of H2{{\rm{H}}_2} is as follows:

σ1s2\sigma 1{s^2}

The molecular orbital configuration of H2{\rm{H}}_{\rm{2}}^ - is as follows:

σ1s2σ1s1\sigma 1{s^2}{\sigma ^*}1{s^1}

Hence, H2{{\rm{H}}_2} is diamagnetic and H2{\rm{H}}_{\rm{2}}^ - is paramagnetic.

Ans:

O2 is paramagnetic, while O22{\rm{O}}_2^{2 - } and O22+{\rm{O}}_2^{2 + } are diamagnetic.

F2 and F22{\rm{F}}_2^{2 - } are diamagnetic and F22+{\rm{F}}_2^{2 + } is paramagnetic.

N2{{\rm{N}}_2} and N22+{\rm{N}}_2^{2 + } are diamagnetic while N2+{\rm{N}}_{\rm{2}}^{\rm{ + }} and N2{\rm{N}}_{\rm{2}}^ - are paramagnetic.

C2{{\rm{C}}_2} and C22{\rm{C}}_{\rm{2}}^{2 - } are diamagnetic while C22+{\rm{C}}_{\rm{2}}^{{\rm{2 + }}} is paramagnetic.

B2{{\rm{B}}_2} is paramagnetic while B22+{\rm{B}}_{\rm{2}}^{{\rm{2 + }}} and B22{\rm{B}}_{\rm{2}}^{2 - } are diamagnetic.

H2{{\rm{H}}_2} is diamagnetic and H2{\rm{H}}_{\rm{2}}^ - is paramagnetic.